OpenStudy (anonymous):

pyramid A is a square pyramid with a base side length of 8 inches and a height of 9 inches. pyramid B has a volume of 2,880 in^3. how many times bigger is the volume of pyramid B than pyramid A? a) 15 b)32 c)54 d)75 medal and fan will be given

OpenStudy (anonymous):

@paki can you help me please

OpenStudy (anonymous):

@campbell_st hi can you help me please

OpenStudy (anonymous):

@iambatman can you help me with this please

OpenStudy (anonymous):

@ganeshie8 hi can you help me

OpenStudy (anonymous):

@thomaster hey can you help me please

OpenStudy (anonymous):

@campbell_st hi can you help me with this please

OpenStudy (campbell_st):

well find the volume of A \[V= \frac{1}{3} \times base.area \times height\] \[V = \frac{1}{3} \times 8 \times 8 \times 9\] find the volume... then compare it to B

OpenStudy (anonymous):

okok doing that now

OpenStudy (anonymous):

i got 192 @campbell_st

OpenStudy (anonymous):

i did my own calculation and i got 93.3 but it isnt one of the answer choices @campbell_st

OpenStudy (campbell_st):

thats the correct volume... 192 to then its 2880/192 = and that's the answer.

OpenStudy (anonymous):

ooo makes sense thank you so much!